OpenStudy (anonymous):

Find two numbers with a sum of 20 and a difference of 14. A. 4 and 18 B. –3 and –17 C. 2 and 16 D. 3 and 17

OpenStudy (anonymous):

let two nos be a b a+b=20 a-b=14 solve them!!!

OpenStudy (callisto):

let x and y be the number respectively x+y = 20 -(1) x-y = 14 (2) (1) +(2) 2x = 34 x=17 Put x=17 into (1) y=20-17 =3 therefore the 2 numbers are 3 and 17

OpenStudy (anonymous):

good job